HUBBLE SPACE TELESCOPE DAILY REPORT #2135 PERIOD COVERED: 0000Z (UTC) 06/01/98 - 0000Z (UTC) 06/02/98 Daily Status Report as of 153/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed NIC/1/2/3 7226 (Search for Massive Jupiters) The Near Infrared Camera (Cameras One, Two and Three) was used to make a corographic search for massive planets around the nearby, young main sequence star RE1147+12. Because of the extreme youth of this object, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed NIC/1/2 7849 (Metallicity and the Cepheid Distance Scale) The Near Infrared Camera (Cameras One and Two) was used to make observations of Cepheid variable stars in the IC4182B galaxy. The local distance scale to galaxies is largely based on the Cepheid period-luminosity relation. Motivated by the goal of measuring an accurate value of the Hubble constant, HST has been used successfully to discover scores of Cepheid variables in each of about 20 galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 7623 (WF/PC-2 Cycle 7 Internal Flats) The WF/PC-2 was used to make a set of internal Cycle 7 calibration flats which are designed to monitor the pixel to pixel flat field response and the visflat lamp degradation, as well as detect any possible changes due to contamination. The calibration observations were completed as planned, and no problems were reported. 1.6 Completed Five Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.7 Completed STIS/CCD/WF/PC-2/NIC/1/2 7569 (The Nuclear Region of NGC-4151) The WF/PC-2, Space Telescope Imaging Spectrograph (CCD) and the Near Infrared Camera (Cameras One and Two) were used to make observations of the nuclear region of Seyfert galaxy NGC-4151 which will be studied in several ways: 1) low-resolution spectra over the entire UV and visible range will be obtained; 2) slitless spectra of strong emission lines at moderate resolution and narrow-band WF/PC-2 images will be used to map the dynamics of the regions; 3) Echelle spectra over the entire UV bandpass will be used to study the intrinsic absorption in the nucleus; and 4) parallel imaging will be performed of the outer galaxy and other galaxies or QSOs. The observations were completed as planned, and no problems were reported. 1.8 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.9 Completed WF/PC-2 7505 (The Hubble Diagram for Distant Supernovae Measuring Cosmic Deceleration and Global Curvature) The WF/PC-2 was used to make observations of Super Nova SN1998E in order to produce a precise Hubble diagram from measuring cosmic deceleration and global curvature. HST can make an essential contribution by providing photometry with substantially higher precision and better temporal coverage than can be obtained from the ground. The observations were completed as planned, and no anomalies were reported. 1.10 Completed Two Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.11 Completed NIC/1/2 7690 (NICMOS Pointed Super Flats) The Near Infrared Camera (Cameras One and Two) was used to make super flat field calibration observations. The observations will be used to build high signal/noise ratio NICMOS flat fields for the different filters. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed STIS/CCD 7639 (STIS CCD Imaging Sensitivity Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the globular cluster NGC-5139 (a.k.a., Omega Cen) in order to monitor the CCD's sensitivity. This proposal should give a check on the long-term, full-field photometric stability of the instrument. A bias observations and a dark calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed NIC/1/2 7864 (The Host Galaxies of Radio-Quiet High- Redshift Quasars) The Near Infrared Camera (Cameras One and Two) was used to make observations of the radio-quite, high red shifted quasar MZZ11408 and its host galaxy. The underlying causes of the rapid cosmological evolution of the quasar population are a mystery around which are woven many different strands of contemporary cosmology. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 20 Successful : 20 2.3 Operations Notes: Due to the MCE-1 reset reported in yesterday's report, a status buffer message was received at 152/1414Z which indicated that the MAMA-1 high voltage had not been allowed to ramp down since an event flag had not been set nominally by the flight software. HSTAR 6501 documents this event. The SSR EDAC error counter was cleared five times using ROP SR-1. The SSR science playback pointer was reset at 153/0217Z per ROP SR-3. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH